# Many-body quantum dimerization in 2D atomic arrays

### v1

@@ Abstract
The study of photon-mediated atom-atom interactions in atomic arrays coupled to different photonic environments has been an extensive field of research in the past years. Two main distinct scenarios have been studied. In the first one, atoms interact via a process of coherent emission and reabsorption leading to an effective coherent Hamiltonian, a topic extensively explored in the context of cavity QED [1]. A similar effect can be achieved when atoms are coupled to a photonic band-gap material. In this case, the evanescent field in the photonic gap can be employed to engineer long-range interactions [2, 3, 4, 5, 6], which can be exploited for quantum simulation of long-range spin models [7, 8, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. The second one considers the coupling of emitters to radiative modes, thus being intrinsically dissipative. This approach began with Dicke’s seminal work [19], where multiple atoms are coupled to a single radiative mode, and predicted the occurrence of collective emission ruled by sub-radiant and super-radiant states.
More recently, significant interest has been directed toward exploring the fate of these collective effects when interactions among emitters acquire spatial dependence, as occurring in ordered atomic arrays coupled to an extended multimode photonic environment. This has been extensively studied in the context of 1D waveguide QED [20, 21], where multiple emitters couple to light confined within a one-dimensional channel, either at optical [22, 23, 24, 25] or microwave frequencies [26, 27, 28, 29]. In this case, light confinement gives rise to strong photon correlations, leading to the emergence of repulsive [30, 31, 32, 33, 34, 35, 36, 37] and attractive [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50] sub-radiant photonic states. While most studies have focused on either the few-excitation regime or the specular scenario of super-radiance [51, 52, 53], where all emitters are initially excited, more complex many-body states are expected to emerge at finite filling fraction. This question was recently addressed in studies showing that, at half-filling of excitations, the most sub-radiant state is composed of a product of dimerized (singlet) states [54, 55, 56], a result closely connected to the steady-state dimerization [57, 58] recently experimentally observed in circuit QED [59].
@@ I Model
We consider an ordered array of $N$ atoms, each with ground and excited states $|g\rangle$ and $|e\rangle$ separated by a distance $d$ and coupled to a photonic bath. The spontaneous emission rate of a single atom from the excited state into the photonic bath is denoted by $\gamma$. As long as the atom-photon dynamics, which is set by $\gamma$, occurs on a time-scale slower than the photon propagation time across the array, the system dynamics is efficiently captured within the Born-Markov approximation. Under this assumption, the photonic degrees of freedom can be adiabatically eliminated, leading to the following Lindblad master equation that governs the atomic system’s dynamics [84, 85] ($\hbar=1$):
@@
where $\hat{\sigma}_{+}^{i}=|e\rangle\langle g|_{i}$ and $\hat{\sigma}_{-}^{i}=|g\rangle\langle e|_{i}$ are raising and lowering spin-1/2 operators that create or destroy an excitation on the $i$-th atom, respectively. In Eq. (1) the anticommutator part of the Lindblad master equation has been included into the non-hermitian Hamiltonian:
$\hat{H}_{\rm eff}=\sum_{i,j}G_{ij}\hat{\sigma}_{+}^{i}\hat{\sigma}_{-}^{j}\,,$ (2)
and we defined the dissipation matrix $\Gamma_{ij}=-{\rm Im}\{G_{ij}\}$. Both objects are determined by the function $G_{ij}$, which encodes the coherent and dissipative long-range photon-mediated interactions among the atoms and it is directly related to the electromagnetic Green’s function of the photonic environment [86, 87]. In this work, we focus our discussion on 2D atomic arrays either coupled to light confined within a 2D photonic waveguide [69, 8, 9, 10, 11] or embedded in free space [86, 87, 88, 89, 90], as sketched in Fig.  1. In the first case, we assume the waveguide displays a quadratic and isotropic dispersion relation. Under this assumption, the photon-mediated atom-atom interactions are governed by [69, 8, 9, 10, 11]:
@@
where $k_{0}=2\pi/\lambda_{0}$ is the photon wavevector, whose corresponding frequency is resonant with the atomic transition of wavelenght $\lambda_{0}$. The term $x_{ij}=|\mathbf{x}_{i}-\mathbf{x}_{j}|$ denotes the distance between atoms in the 2D array, and $\mathcal{J}_{0}$ and $\mathcal{Y}_{0}$ represent the zeroth-order Bessel functions of the first and second kinds, respectively. In the second case, we assume that the atoms have orthogonal polarization with respect to the array plane. In this case, the correlated free-space emission is given by [86]:
$G^{2D\rm,free}_{ij}=\frac{3\gamma_{0}}{(4(k_{0}x_{ij})^{3}}e^{ik_{0}x_{ij}}\Big{(}1-ik_{0}x_{ij}-(k_{0}x_{ij})^{2}\Big{)}.$ (4)
Finally, for comparison with previous results, we also consider the case of multiple atoms coupled to a one-dimensional waveguide [20]. In this case, the effective interactions are described by the function [20]:
@@
Figure 2: (a) Singlet and triplet decay rates in unit of $\gamma$ for two atoms as function of the inter-atomic distance for the three models of Eqs. (3)-(5). The shaded orange area indicate the inter-atomic distances where a NN dimer is expected to be less radiant than long range dimers and triplet states. (b) Fidelity as function of the inter-atomic distance between the dimer covering ansatz and the least radiant state within the $N/2$ subspace of the effective Hamiltonian (2) for the three considered models. Here we considered a $N=4,8,16$ chain for the 1D waveguide and a $N=2\times 2,4\times 2,4\times 4$ square lattice for the other two cases. The color scale indicates the decay rate $\gamma_{s=1}$ of the least radiant state. The results are obtained via exact diagonalization of the non-Hermitian Hamiltonian $\hat{H}_{\text{eff}}$.
@@ II.1 Dimerization in 1D waveguide QED
In this section, we build some physical intuition for identifying the least radiant state of the system at half-filling. For an atomic array coupled to a 1D waveguide, it is well established that subradiant states in the two-excitation sector are typically composed of repulsive (or so-called fermionic) excitations [86, 30, 31, 32, 33, 34, 35, 36, 37]. For these states, the two particles are maximally far from each other and from the edge of the system as well, thus reducing the likelihood of photons scattering out of the system. At higher fillings, the mechanism is similar by shrinking the multiparticle distribution accordingly [54]. Instead, half-filling is the limiting case where excitations can spread the least and one might intuitively expect the least radiant state to emerge from each excitation being shared between, in principle, any pair of atoms. This intuition was formalized for 1D waveguide QED systems in Ref. [54], which demonstrated that the least radiant state at half-filling is represented by a product state of all nearest-neighbor (NN) dimers (singlet states):
$|\phi_{D}\rangle=\prod_{j\in\rm odd}\hat{D}_{j,j+1}|0\rangle\,.$ (6)
@@
is the dimer creation operator. The formation of dimers only between nearest-neighbor atoms can be understood through the following argument. In the Dicke model limit of Eq. (5), the dark subspace with zero decay rate consists of all possible dimers, including both short- and long-range configurations. As one moves smoothly away from this limit, the spatial dependence of $G_{ij}^{1D}$ induces a spatial dependence of the collective decay rates as well. As we will see below, short-range dimer coverings are favoured over long-range ones, as these configurations minimize the decay rate of pairs of emitters.
@@
If this condition is not met, the formation of triplets or long-range singlet pairings will result in a state with a lower decay rate. This can occur in the 1D waveguide scenario, where the condition given in Eq. (9) needs to be fine tuned for large arrays. Indeed, since the spatial periodic dependence of the decay rate (see Fig. 2(a)) is generally incommensurate with the lattice spacing, there can be an atom-atom distance at which the decay of long-range dimers or triplet states becomes arbitrarily smaller than that of nearest-neighbor ones (see App. A). At such distances, the least radiant state will start incorporating long-range dimers or triplet states, reducing its resemblance to the dimerized ansatz given in Eq. (6).
This can be seen in the the first panel of Fig. 2(b), where we compute the global fidelity, $F=|\langle\phi_{D}|\psi_{\rm s=1}\rangle|^{2}$, between the least radiant state of the $N/2$ excitation subspace, $|\psi_{\rm s=1}\rangle$, where the index $s=1$ is used to order states in decreasing value of decay rate, and the dimer product state given in Eq. (6). The fidelity rapidly decreases with interatomic distance as the array size increases, due to the argument explained above.
@@ II.2 Dimerization in 2D arrays
In two spatial dimensions, both for the 2D WQED and free-space cases, repulsive two-excitation sub-radiant states have been identified [69]. These findings suggest that a pairwise entanglement structure at half-filling fraction could also emerge in this context. The condition given in Eq. (9) is satisfied within the distance range highlighted by the orange-shaded area in Fig. 2(a). This region corresponds to where the NN dimer decay rate remains lower than the decay associated with long-range dimer pairs and the triplet state. This behavior arises due to the dampening of the collective decay rates, which scale as $1/\sqrt{k_{0}d}$ in a 2D waveguide and as $1/(k_{0}d)$ in free space. This suggests that nearest-neighbor dimers should be favored in forming the least radiant state. However, in contrast with the 1D waveguide case, in this scenario there are different possible dimer coverings. In the following discussion, we focus on the non-frustrated case of a square lattice consisting of $N=N_{x}\times N_{y}$ atoms. For this scenario, we conjecture that the least radiant state at half-filling is an equal coherent superposition of all nearest-neighbor quantum dimers, as illustrated in Fig. 1(c). We thus propose the following ansatz:
$|\phi_{\rm RVB}\rangle=\sum_{\alpha}\prod_{i,j\in\mathcal{D}_{\alpha}}\hat{D}_{i,j}|0\rangle\,,$ (10)
@@
where the sum runs over all possible NN dimer coverings $\mathcal{D}_{\alpha}=\{(i_{1},i_{2}),(i_{3},i_{4})....\}_{\alpha}$ and the pair of indices $(i_{a},i_{b})$ labels the position of the two atoms forming a given dimer. For instance, for a $N=2\times 2$ lattice labelling the atoms in the first raw as $1,2$ and in the second as $3,4$ there are two possible dimer coverings given by $\mathcal{D}_{1}=\{(1,2),(3,4)\}_{1}$ and $\mathcal{D}_{2}=\{(1,3),(2,4)\}_{2}$. As the size of the array increases, the number of possible dimer coverings grows exponentially, but all the possible dimer coverings in the lattice can be found numerically (see App. B). Note that this ansatz coincides with the “Resonating Valence bond” (RVB) liquid phase wavefunction for a square lattice, as proposed by Anderson [72, 71] and also appearing in the Rokhsar-Kivelson (RK) quantum dimer model at the so-called RK point [73, 74, 75]. To probe whether our conjecture holds for a 2D atomic array, we compute the global fidelity $F=|\langle\phi_{\rm RVB}|\psi_{\rm s=1}\rangle|^{2}$, between the least radiant state at half-filling and the RVB ansatz. The global fidelity as a function of inter-atomic distance is shown in Fig. 2(b) for the three cases described in Eqs. (3)-(5) and different system sizes. In all cases, the fidelity decreases as the system size increases. Interestingly, the ansatz appears to be more robust for 2D arrays than the 1D scenario, with the maximum fidelity extending over a larger range of finite distances in the region where Eq. (9) holds. Specifically, the fidelity maximum remains relatively flat at shorter distances in the 2D waveguide case. In free space, instead, the maximum in fidelity is achieved at an intermediate distance $k_{0}d\sim 0.4\pi$. This effect is most likely due to the fact that strong near-field coherent interactions occur at short distances, thus shifting the optimal regime to larger inter-atomic distances.
@@
Figure 3: (a) Global Fidelity, $F$, and (b) global Infidelity density, $\bar{I}_{F}$, between the RVB ansatz and the least radiant state for the models (3)-(5) as function of the system size for two different lattice geometries, as indicated. For the 1D and 2D waveguide case we fixed $k_{0}d=0.1\pi$ while for the free space case we fixed the inter-atomic distance to $k_{0}d=0.42\pi$.

### v4

@@ Abstract
The study of photon-mediated atom-atom interactions in atomic arrays coupled to different photonic environments has been an extensive field of research in the past years. Two main distinct scenarios have been studied. In the first one, atoms interact via a process of coherent emission and reabsorption leading to an effective coherent Hamiltonian, a topic extensively explored in the context of cavity QED [1]. A similar effect can be achieved when atoms are coupled to a photonic band-gap material. In this case, the evanescent field in the photonic gap can be employed to engineer long-range interactions [2, 3, 4, 5, 6], which can be exploited for quantum simulation of long-range spin models [7, 8, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. The second one considers the coupling of emitters to radiative modes, thus being intrinsically dissipative. This approach began with Dicke’s seminal work [19], where multiple atoms located at the same position are coupled to the electromagnetic field, and predicted the occurrence of collective emission ruled by sub-radiant and super-radiant states.
More recently, significant interest has been directed toward exploring the fate of these collective effects when interactions among emitters acquire spatial dependence, as occurring in ordered atomic arrays coupled to an extended multimode photonic environment. This has been extensively studied in the context of 1D waveguide quantum electrodynamics (WQED)  [20, 21], where multiple emitters couple to light confined within a one-dimensional channel, either at optical [22, 23, 24, 25] or microwave frequencies [26, 27, 28, 29]. In this case, light confinement gives rise to strong photon correlations, leading to the emergence of repulsive [30, 31, 32, 33, 34, 35, 36, 37] and attractive [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50] sub-radiant photonic states. While most studies have focused on either the few-excitation regime or the specular scenario of super-radiance [51, 52, 53], where all emitters are initially excited, more complex many-body states are expected to emerge at finite filling fraction. This question was recently addressed in studies showing that, at half-filling of excitations, the most sub-radiant state is composed of a product of dimerized (singlet) states [54, 55, 56], a result closely connected to the steady-state dimerization [57, 58] recently experimentally observed in circuit QED [59].
@@ I Model
We consider an ordered array with lattice constant $d$ made of $N$ two-level atoms, each with ground and excited states $|g\rangle$ and $|e\rangle$, separated by a transition frequency $\omega_{eg}$. The atoms are coupled at position $\mathbf{x_{i}}$ to a generic photonic bath described by a collection of photonic modes indexed by $n$ and having frequencies $\omega_{n}$. The Hamiltonian that describes this coupled atom-light interacting system under the rotating-wave and dipole approximations reads ($\hbar=1$):
$\begin{split}&\hat{H}=\hat{H}_{0}+\hat{H}_{\rm int}\,,\\ &\hat{H}_{0}=\sum_{i}\omega_{eg}\hat{\sigma}_{+}^{i}\hat{\sigma}_{-}^{i}+\sum_{n}\omega_{n}\hat{a}_{n}^{\dagger}\hat{a}_{n}\,,\\ &\hat{H}_{\rm int}=\sum_{i}\sum_{n}\left[g_{n}(\mathbf{x_{i}})\hat{a}_{n}\hat{\sigma}_{+}^{i}+\text{H.c.}\right]\,,\end{split}$ (1)
where the index $i$ runs over all atoms in the array, and $\hat{\sigma}_{+}^{i}=(\hat{\sigma}_{-}^{i})^{\dagger}=|e\rangle_{i}\langle g|$ and $\hat{a}_{n}^{\dagger}$ denote the atomic and bosonic creation operators, respectively. The interaction between the atoms and the photonic bath is ruled by the atom-photon coupling constant
$g_{n}(\mathbf{x}_{i})=\sqrt{\frac{\omega_{n}}{2\epsilon_{0}}}\mathbf{d}^{i}_{eg}\cdot\mathbf{\Phi}_{n}(\mathbf{x}_{i})\,,$ (2)
with $\epsilon_{0}$ being the vacuum permittivity, $\mathbf{d}^{i}_{eg}=\langle e|\hat{\mathbf{d}}^{i}|g\rangle$ the dipole moment of the i-th atom and $\mathbf{\Phi}_{n}(\mathbf{x}_{i})$ the spatial mode eigenfunction of the field normalized according:
[…]
These functions are solutions of the Helmholtz equation, determined by the boundary conditions and, consequently, by the geometry of the system. In the next sections, we examine how this influences the photon-mediated interactions between atoms in various geometries, within the framework of the Born-Markov approximation.
I.1 Effective spin model
To simplify the description of this complex interacting system, we assume that the standard conditions for applying the Born-Markov approximation are fulfilled. These conditions are satisfied when the density of states of the photonic bath is sufficiently smooth and when time-retardation effects—arising from the finite speed of light $c$—can be neglected. The latter can be formalized by requiring that the timescale of atom-photon dynamics, governed by the spontaneous emission rate $\gamma$ of a single atom into the photonic bath, is much longer than the time it takes for a photon to propagate across the system, i.e., $\gamma\ll c/(Nd)$ In this regime, light-mediated interactions between atoms can be treated as effectively instantaneous. Equivalently, the strongly dispersive nature of the atoms ensures that the resulting light-matter dressed excitations are predominantly atomic in character. Under these assumptions, the photonic degrees of freedom can be adiabatically eliminated, yielding the following Lindblad master equation that governs the dynamics of the atomic system [84, 85]:
$\dot{\hat{\rho}}(t)=-i[\hat{H}_{S},\hat{\rho}(t)]+\sum_{ij}\Gamma_{ij}\left(2\hat{\sigma}_{-}^{j}\hat{\rho}\hat{\sigma}_{+}^{i}-\{\hat{\sigma}_{+}^{i}\hat{\sigma}_{-}^{j},\hat{\rho}\}\right)\,,$ (4)
where
$\hat{H}_{S}=\sum_{i}\omega_{eg}\hat{\sigma}_{+}^{i}\hat{\sigma}_{-}^{i}+\sum_{i,j}{\rm Re}\{G_{ij}\}\hat{\sigma}_{+}^{i}\hat{\sigma}_{-}^{j}\,,$ (5)
describes the coherent collective dynamics of the atomic system while the matrix $\Gamma_{ij}=-{\rm Im}\{G_{ij}\}$ encodes the collective dissipation. Coherent and dissipative long-range photon-mediated interactions among the atoms are ruled by the function $G_{ij}$, which is directly related to the electromagnetic Green’s function of the photonic environment [86, 87] and is given by:
$G_{ij}=\,\lim_{s\to 0^{+}}\sum_{n}\frac{g_{n}(\mathbf{x_{i}})^{*}g_{n}(\mathbf{x_{j}})}{is-(\omega_{a}-\omega_{n})}.$ (6)
It is convenient to rewrite the Lindblad master equation written in Eq. (4) in the form
@@ I.1 Effective spin model
where the anticommutator part of the master equation has been incorporated into an effective spin model non-Hermitian Hamiltonian, defined as:
$\hat{H}_{\rm eff}=\sum_{i,j}G_{ij}\hat{\sigma}_{+}^{i}\hat{\sigma}_{-}^{j}\,.$ (8)
This effective Hamiltonian inherently encodes photon-photon interactions, as can be seen by recasting it as a hard-core boson model (see Ref. [20] for the 1D waveguide case and Ref. [69] for the 2D arrays). In the following, we will explore various scenarios characterized by different forms of the interaction matrix $G_{ij}$.
I.2 Interaction matrix
In this work, we focus our discussion on the following scenarios. The first two are sketched in Fig. 1.
  • In the first case, we consider a 2D atomic array coupled to light confined within a 2D photonic waveguide [69, 8, 9, 10, 11]. We assume the waveguide displays a quadratic and isotropic continuous dispersion relation $\omega_{\mathbf{k}}=\omega_{c}+A|\mathbf{k}|^{2}$, where $\omega_{\mathbf{k}}$ are the field frequencies labeled by the wavevector $\mathbf{k}$, $\omega_{c}$ is the dispersion cutoff and we consider $A>0$. Under this assumption, the rate at which an atom emits a photon into the structure is given by $\gamma=|g_{\mathbf{k}_{0}}(\mathbf{x_{i}})|^{2}/(4A)$ and the photon-mediated atom-atom interaction matrix reads [69, 8]:
@@ I.2 Interaction matrix
    where $k_{0}=2\pi/\lambda_{0}$ is the photon wavevector, whose corresponding frequency is resonant with the atomic transition of wavelenght $\lambda_{0}$. The term $x_{ij}=|\mathbf{x}_{i}-\mathbf{x}_{j}|$ denotes the distance between atoms in the 2D array, and $\mathcal{J}_{0}$ and $\mathcal{Y}_{0}$ represent the zeroth-order Bessel functions of the first and second kinds, respectively.
  • In the second scenario, we consider a 2D atomic array embedded in free space [88, 89, 90, 91, 92, 93]. In this case, atomic emission is governed by the free-space spontaneous decay rate, $\gamma=d_{eg}^{2}k_{0}^{3}/(3\pi\epsilon_{0})$. Assuming that the atomic dipoles are polarized orthogonally to the array plane, the correlated free-space emission is given by [86, 87]:
    $G^{2D,\mathrm{free}}_{ij}=\frac{3\gamma\,e^{ik_{0}x_{ij}}}{4(k_{0}x_{ij})^{3}}[1-ik_{0}x_{ij}-(k_{0}x_{ij})^{2}]$ (10)
  • Finally, for comparison with previous results, we consider the case of multiple atoms coupled to a one-dimensional waveguide [20]. In this case, assuming a linear dispersion relation for the photons of the form $\omega_{k}=c|k|$, the decay rate of the atom into the waveguide is given by $\gamma=|g_{k_{0}}(x_{i})|^{2}/(2c)$ and the effective interactions are described by the interaction matrix [20]:
@@ I.3 Half-filling subspace
From the exact diagonalization of the half-filling sector, we obtain a set of eigenvalues of the form
$E^{(N/2)}_{s}=\epsilon^{(N/2)}_{s}-i\gamma^{(N/2)}_{s}/2,$ (12)
where the index $s$ orders the states by increasing decay rate. Here, $\epsilon^{(N/2)}_{s}$ and $\gamma^{(N/2)}_{s}$ denote the energy and collective decay rate of the $s$-th state, respectively.
Fixing the excitation sector also allows us to apply exact diagonalization methods within a reduced Hilbert space of dimension given by the binomial coefficient $\chi=\binom{N}{N/2}$. This reduction enables us to access system sizes of up to $N=24$ atoms.
Figure 2: (a) Singlet and triplet decay rates in unit of $\gamma$ for two atoms as function of the inter-atomic distance for the three models of Eqs. (9)-(11). The shaded orange area indicate the inter-atomic distances where a NN dimer is expected to be less radiant than long range dimers and triplet states. (b) Fidelity as function of the inter-atomic distance between the dimer covering ansatz and the least radiant state within the $N/2$ subspace of the effective Hamiltonian (8) for the three considered models. Note that we use Eq. (13) as the ansatz for the 1D waveguide case, and Eq. (17) for the two-dimensional setting. We considered a $N=4,8,16$ chain for the 1D waveguide and a $N=2\times 2,4\times 2,4\times 4$ square lattice for the other two cases. The color scale indicates the decay rate $\gamma_{s=1}$ of the least radiant state. The results are obtained via exact diagonalization of the non-Hermitian Hamiltonian $\hat{H}_{\text{eff}}$.
@@ II Many-body quantum dimerization
In this section, we discuss the occurrence of many-body dimerization across the different scenarios introduced in Sec. I.2. We begin in Sec. II.1 with the 1D waveguide case, in order to build physical intuition for identifying the least radiant state of the system at half-filling as a product of dimer states. We then examine in Sec. II.2 how this mechanism can lead to the emergence of an RVB state as the least radiant state in a 2D array. Finally, in Sec. II.3, we characterize this state in detail.
@@ II.1 Dimerization in 1D waveguide QED
For an atomic array coupled to a 1D waveguide, it is well established that subradiant states in the two-excitation sector are typically composed of repulsive (or so-called fermionic) excitations [86, 30, 31, 32, 33, 34, 35, 36, 37]. For these states, the two particles are maximally far from each other and from the edge of the system as well, thus reducing the likelihood of photons scattering out of the system. This becomes particularly clear at the two-excitation level, by solving the two-body problem described by Hamiltonian (8) in the relative coordinate frame [20, 69]. In this framework, the hard-core photon-photon interactions are mapped onto an impurity term, which captures the atomic saturation, i.e., the inability of an atom to absorb more than one excitation. The repulsive behavior of the excitations thus emerges from the effective hard-wall condition imposed by this impurity potential.
At higher fillings, the mechanism is similar by shrinking the multiparticle distribution accordingly [54]. Instead, half-filling is the limiting case where excitations can spread the least and one might intuitively expect the least radiant state to emerge from each excitation being shared between, in principle, any pair of atoms. This intuition was formalized for 1D waveguide QED systems in Ref. [54], which demonstrated that the least radiant state at half-filling is represented by a product state of all nearest-neighbor (NN) dimers (singlet states):
$|\phi_{D}\rangle=\prod_{j\in\rm odd}\hat{D}^{\dagger}_{j,j+1}|0\rangle\,.$ (13)
@@
is the dimer creation operator. The formation of dimers only between nearest-neighbor atoms can be understood through the following argument. In the Dicke model limit of Eq. (11), the dark subspace with zero decay rate consists of all possible dimers, including both short- and long-range configurations. As one moves smoothly away from this limit, the spatial dependence of $G_{ij}^{1D}$ induces a spatial dependence of the collective decay rates as well. As we will see below, short-range dimer coverings are favoured over long-range ones, as they minimize the decay rates of emitter pairs. Consequently, a state predominantly composed of nearest-neighbor dimers is filtered out from the degenerate singlet subspace.
@@
If this condition is not met, the formation of triplets or long-range singlet pairings will result in a state with a lower decay rate. In the 1D waveguide scenario the condition given in Eq. (16) needs to be fine tuned for sufficiently large arrays. Indeed, since the spatial periodic dependence of the decay rate (see Fig. 2(a)) is generally incommensurate with the lattice spacing, there can be an atom-atom distance at which the decay of long-range dimers or triplet states becomes arbitrarily smaller than that of nearest-neighbor ones (see App. A). At such distances, the least radiant state will start incorporating long-range dimers or triplet states, reducing its resemblance to the dimerized ansatz given in Eq. (13).
This can be seen in the the first panel of Fig. 2(b), where we compute the global fidelity, $F=|\langle\phi_{D}|\psi_{\rm s=1}\rangle|^{2}$, between the least radiant state of the $N/2$ excitation subspace, $|\psi_{\rm s=1}\rangle$, labeled by the index $s=1$ defined in Eq. (12), and the dimer product state given in Eq. (13). The fidelity rapidly decreases with interatomic distance as the array size increases, due to the argument explained above.
@@ II.2 Dimerization in 2D arrays
In two spatial dimensions, both for the 2D WQED and free-space cases, repulsive two-excitation sub-radiant states have been identified [69]. These findings suggest that a pairwise entanglement structure at half-filling fraction could also emerge in this context. The condition given in Eq. (16) is satisfied within the distance range highlighted by the orange-shaded area in Fig. 2(a). This region corresponds to where the NN dimer decay rate remains lower than the decay associated with long-range dimer pairs and the triplet state. This behavior arises due to the dampening of the collective decay rates, which scale as $1/\sqrt{k_{0}d}$ in a 2D waveguide and as $1/(k_{0}d)$ in free space. This suggests that nearest-neighbor dimers should be favored in forming the least radiant state. However, in contrast with the 1D waveguide case, in this scenario there are different possible dimer coverings. In the following discussion, we focus on the non-frustrated case of a square lattice consisting of $N=N_{x}\times N_{y}$ atoms. Given that it is reasonable to assume no particular dimer covering is dissipatively favored over another, we conjecture that, in this scenario, the least radiant state at half-filling is a coherent and equal superposition of all nearest-neighbor quantum dimers, as illustrated in Fig. 1(c). We thus propose the following ansatz:
$|\phi_{\rm RVB}\rangle=\sum_{\alpha}\prod_{i,j\in\mathcal{D}_{\alpha}}\hat{D}^{\dagger}_{i,j}|0\rangle\,,$ (17)
@@
where the sum runs over all possible NN dimer coverings $\mathcal{D}_{\alpha}=\{(i_{1},i_{2}),(i_{3},i_{4})....\}_{\alpha}$ and the pair of indices $(i_{a},i_{b})$ labels the position of the two atoms forming a given dimer. For instance, for a $N=2\times 2$ lattice labeling the atoms in the first raw as 1,21,21 , 2 and in the second as 3,43,43 , 4 there are two possible dimer coverings given by $\mathcal{D}_{1}=\{(1,2),(3,4)\}_{1}$ and $\mathcal{D}_{2}=\{(1,3),(2,4)\}_{2}$. As the size of the array increases, the number of possible dimer coverings grows exponentially, but all the possible dimer coverings in the lattice can be found numerically (see App. B). Note that this ansatz coincides with the “Resonating Valence bond” (RVB) liquid phase wavefunction for a square lattice, as proposed by Anderson [72, 71] and also appearing in the Rokhsar-Kivelson (RK) quantum dimer model at the so-called RK point [73, 74, 75].
To probe whether our conjecture holds for a 2D atomic array, we compute the global fidelity $F=|\langle\phi_{\rm RVB}|\psi_{\rm s=1}\rangle|^{2}$, between the least radiant state at half-filling, again labeled by the index $s=1$ defined in Eq. (12), and the RVB ansatz. The global fidelity as a function of inter-atomic distance is shown in Fig. 2(b) for the three cases described in Eqs. (9)-(11) and different system sizes. In all cases, the fidelity decreases as the system size increases. This drop in fidelity can be attributed to the fact that both dissipative and coherent interactions induce mixing among the various dimer configurations that span the original singlet subspace, thereby perturbing the RVB-like state. Interestingly, the ansatz appears to be more robust for 2D arrays than the 1D scenario, with the maximum fidelity extending over a larger range of finite distances in the region where Eq. (16) holds. Specifically, the fidelity maximum remains relatively flat at shorter distances in the 2D waveguide case. In free space, instead, the maximum in fidelity is achieved at an intermediate distance $k_{0}d\sim 0.4\pi$. This effect is most likely due to the fact that strong near-field coherent interactions occur at short distances, thus shifting the optimal regime to larger inter-atomic distances.
@@
Figure 3: (a) Global Fidelity, $F$, and (b) global Infidelity density, $\bar{I}_{F}$, between the dimer covering ansatz (Eq. (13) for the 1D WQED case and Eq. (17) for the 2D case) and the least radiant state for the models (9)-(11) as function of the system size for two different lattice geometries, as indicated. For the 1D and 2D waveguide case we fixed $k_{0}d=0.1\pi$ while for the free space case we fixed the inter-atomic distance to $k_{0}d=0.42\pi$.